OpenStudy (anonymous):

If a = -12/5, then 5x + 2y = 6 and 3x - ay = 4 are parallel. i got true?

OpenStudy (anonymous):

I don't believe they are parallel.

OpenStudy (anonymous):

i see now

OpenStudy (anonymous):

\[3x - ay = 4 \implies 3x - \frac{-12}{5}y = 4\]\[\implies 5x + 4y = 4\] Which is not the same slope as 5x + 2y = 6